A toy car may be released from rest at a position on an inclined track. It is known to have an efficiency of 52% in its ability to transform its available GPE into KE at the bottom of the track. How much GPE should the car be given so that it has 0.87 Joules of KE at the bottom of the track?


1
Expert's answer
2021-02-25T11:25:20-0500
"Efficiency=\dfrac{KE\ at\ the\ bottom\ of\ track}{Initial\ GPE},""Initial\ GPE=\dfrac{KE\ at\ the\ bottom\ of\ track}{Efficiency},""Initial\ GPE=\dfrac{0.87\ J}{0.52}=1.673\ J."